The paint is 21 years newer that the brush. The paint is also one fourth as old as the brush. How old is each one?
Answer by ptaylor(2198)

Let x=brush's age
then x-21=paint's age
So our equation to solve is:
x-21=(1/4)x multiply each term by 4
4x-84=x subtract 4x from each side
4x-4x-84=-4x+x collect like terms
-3x=-84
x=28 years old---age of the brush
x-21=28-21=7 years old---age of the paint
CK
28-7=21
21=21---ok
(1/4)*28=7
7=7---ok
Hope this helps------ptaylor